[figure: 9.7 cm]

MAGNETIZATION AND MAGNETIC ORDERING IN QUASI

LOW-DIMENSIONAL MAGNET CsDy(MoO_{4})_{2}

E.N. Khatsko^{a}, A.S. Cherny^{a} and A.N. Knigavko^{b}

aInstitute for Low Temperature Physics, National Academy of Sciences of Ukraine,

47 Lenin Ave., 310064 Kharkov, Ukraine bNational Chiao Tung University, Hsinhchu, Taiwan 30050, R.O.C.

Magnetization of single crystals of CsDy(MoO_{4})_{2} has been investigated in the temperature range 0.6-4.2 K in magnetic fields up to 20 kOe. We have measured the field dependence of magnetization both in magnetically ordered (T_{N}=1.36 K) and in paramagnetic states for an external magnetic field applied along the principal crystal axes. c-direction demonstrates a very weak linear M(H) dependence. This is the reason to conclude that the ground state magnetic structure at H = 0 are lying in a − b plane. The magnetization curves in b direction below 1 K demonstrate a steep growth which is similar to ferromagnetic behavior, but with no noticeable remanent magnetization along b directions. The magnetization curves in ordered state along a direction demonstrate some kind of spin reorientation from the initially antiferromagnetic configuration much more clearly than curves along b direction. On the basis of these results and newly established space symmetry of the compound we have performed symmetry analysis of the recent neutron diffraction data [1] and pointed out possible equilibrium magnetic configurations.

[1].W.T. Klooster, A. Zheludev, R.S. Srivastava (to be published)

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : oral

Corresponding author : E.N.Khatsko

Address for correspondence : Institute for Low Temperature Physics 47 Lenin Ave. 310164 Kharkov Ukraine

Email address : khatsko@ilt.kharkov.ua